OpenStudy (anonymous):

square root of 8 x plus 1 equals 5

OpenStudy (anonymous):

show all the work please!

OpenStudy (anonymous):

\[\sqrt{8x +1} = 5\] or \[\sqrt{8x} +1=5\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[(8x +1)^{1/2} =5\] \[8x + 1 =5^2\] \[x = (25 - 1)/8\] x = 3